• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Gilman

    Log entry time 16:08:56 on June 5,2000

    Entry number 43166

    keyword=running shift summary

    beam continuing for first part of shift...
    stops at 8:40 for maintenance

    Target has been put to sleep:

    25%, 60 Hz --> 6%, 14.4 Hz

    ~38% JTvalve, 16g/s flow --> 14% 4.7 g/s flow

    flow decreased over many minutes so cryo better not complain

    JP comes by to look at ice on target. Mike Seely coimes by and is going
    to disconnect lifting controls.

    by end of shift with ~very little coolant the target has warmed to 19.15 degr,
    while the warm return gas gone from ~126 to 147